Circular No. 5730

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

1993 FW J. Luu, University of California at Berkeley; and D. Jewitt, University

of Hawaii, report the discovery of a second very faint object with very slow (3′′/hr) retrograde opposition motion. The object was detected in CCD images obtained with the University of Hawaii’s 2.2-m telescope at Mauna Kea. The object appears stellar in 0′′.8 seeing, with an apparent Mould magnitude R = 22.8 ± 0.1 measured in a 1′′.6-radius aperture. The color V − R = +0.4 ± 0.1.

1993 UT  α_{2000}  δ_{2000}
Mar. 28.41684  12^{h}27^{m}06.^{s}31  −3^{o}00′12′′.7
28.59830  12 27 05.52  −3 00 07.0
29.41934  12 27 01.90  −2 59 42.2
29.42846  12 27 01.85  −2 59 41.8

Computations by B. G. Marsden indicate that 1993 FW is currently between 38 and 56 AU from the earth. Similarity in motion and brightness to 1992 QB_{1} suggests that 1993 FW is another Kuiper Belt candidate, and the orbit selected below is again the direct circle. The object’s phase angle reached a minimum of 0^{o}.002 on Mar. 28.0 UT.

Epoch = 1993 Mar. 14.0 TT  u = 359.^{o}468
Ω = 187.896 }2000.0
a  = 42.451 AU  i = 8.029
1993 TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  V
Mar. 24 12^{h}27.^{m}43 − 3^{o}02.^{′}4 41.456 42.451 176.^{o}0 0.^{o}1 23.2
Apr. 3 12 26.70 − 2 57.4 41.457 42.451 174.0 0.1 23.2
13 12 25.98 − 2 52.4 41.487 42.451 163.9 0.4 23.2
23 12 25.30 − 2 47.7 41.545 42.451 154.0 0.6 23.2
May 3 12 24.70 − 2 43.5 41.631 42.451 144.1 0.8 23.3

COMET SHOEMAKER-LEVY (1993e) Luu and Jewitt also report that their CCD imaging of this comet

with the 2.2-m telescope at Mauna Kea on Mar. 27 showed as many as 17 separate sub-nuclei “strung out like pearls on a string” over a range of 50′′ along p.a. 77^{o}-257^{o}. The brightest sub-nuclei were four or five in from the southwestern end. The appearance was unchanged on Mar. 28.

1993 March 29 Brian G. Marsden